# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

### v3

@@ 4.1 Zero-shot Music Editing Results
AudioLDM 2 23.86 0.21 10.36 0.58 73.48 70.12
MusicGen 23.39 0.21 6.63 0.59 75.46 71.28
SDEdit 25.87 0.22 12.18 0.40 69.38 66.23
DDIM Inversion 22.52 0.21 9.51 0.49 73.10 3.38
MusicMagus 16.23 0.19 5.15 0.55 75.12 74.34
DDPM-Friendly 18.30 0.19 5.16 0.54 75.27 73.86
MEDIC 11.97 0.15 2.49 0.61 79.81 77.29
Table 1: Comparison with baselines on ZoME-Bench with fixed length about structure, content preservation, and CLAP similarity.
We present a comparative study of our Disentangled Inversion Control (DIC) against several established music generation and editing baselines, including AudioLDM 2 (Liu et al., 2023c), MusicGen (Copet et al., 2023), SDEdit (Liu et al., 2023a), DDIM Inversion Ho et al. (2020), MusicMagus (Zhang et al., 2024), and DDPM-Friendly (Manor & Michaeli, 2024). Utilizing the ZoMo-Bench test set we developed, we assess the quality of the generated audio samples, focusing on two key aspects: Content Preservation and Editing Fidelity. The results, summarized in Table 1, lead to the following conclusions: (1) Our method, MEDIC, significantly outperforms both generation-based and inversion-based models in terms of editing fidelity and audio similarity, demonstrating its effectiveness in addressing complex editing tasks. (2) Although DDPM-Friendly and MusicMagus show improvements in essential content preservation, they struggle to maintain high text-audio alignment compared to generation models. MEDIC effectively addresses accumulated errors, achieving high audio similarity and an impressive CLAP score, reflecting its superior editing fidelity. (3) MEDIC also excels in subjective metrics, achieving the best performance in audio quality and content preservation.
Variable Length Comparisons We further evaluate MEDIC against baseline methods in a variable length setting, with the results presented in Table 2. Our analysis leads to the following conclusions: (1) MEDIC outperforms all baselines across all metrics, demonstrating the effectiveness of our approach in variable-length scenarios. (2) Inversion-based baselines experience a significant degradation in CLAP score, while MEDIC maintains the highest CLAP score of 0.51 on the MusicDelta dataset.
Method Objective Metrics Subjective Metrics
@@
AudioLDM 2 24.40 0.22 7.07 0.44 66.37 64.28
MusicGen 27.71 0.23 7.70 0.46 67.41 63.76
SDEdit 28.12 0.24 13.21 0.24 62.80 62.18
DDIM Inversion 23.5 0.21 10.12 0.27 65.94 65.73
MusicMagus 25.6 0.22 7.13 0.43 67.45 67.12
DDPM-Friendly 21.53 0.23 6.68 0.30 66.34 67.28
MEDIC 19.5 0.20 6.58 0.51 71.62 70.18
Table 2: Comparison with baselines in variable length setting.
(a) Spider chart for CLAP score comparisons .
(b) Spider chart for FAD score comparisons.
Figure 3: A comprehensive performance evaluation on the ZoME-bench. We present spider charts of CLAP scores and FAD scores across 10 editing tasks for DDPM-Friendly, DDIM Inversion, MusicMagus, and Disentangled Inversion Control methods.
Fine-grained Comparisons on ZoME-Bench To substantiate the robustness of our method, we delve into fine-grained comparisons across different types of editing. The FAD and CLAP scores are depicted in Figure 3. The insights gleaned from this analysis are as follows: (1) Across all editing types, our Disentangled Inversion Control surpasses other methods, demonstrating its prowess in handling both rigid and non-rigid editing tasks. (2) While baselines exhibit capabilities in handling certain rigid editing scenarios, they fall short in executing non-rigid manipulations, as reflected in their inferior performance, especially in the “Change Genre" and “Change Melody" non-rigid editing types.
4.2 Ablation Study
The ablation studies presented in this section aim to validate the contributions of the Harmonized Attention Control and Disentangled Inversion to our framework’s overall performance. Additionally, we examine the impact of the classifier-free guidance scale on editing outcomes, with a comprehensive analysis included in Appendix D.
4.2.1 Ablation on Attention Control Methods
@@
HAC 11.97 0.15 2.49 0.61
w/o MSA Control 14.13 0.19 2.75 0.56
w/o CA Control 13.78 0.18 2.50 0.58
w/o Harmonic Branch 12.75 0.16 2.59 0.59
Table 3: Ablation results about different control methods. MSA: Mutual Self-Attention, CA: Cross Attention.
To validate the impact of our attention control mechanisms, we perform ablations on the following configurations: Remove Mutual Self Attention Control (w/o MSA Control), Remove Cross Attention Control (w/o CA Control), and without Harmonic Branch (w/o Harmonic Branch). The findings, detailed in Table 3, reveal that: (1) Both cross-attention and mutual self-attention controls individually enhance editing performance. (2) Although the naive combination of mutual self-attention control and cross-attention control improves the preservation and fidelity, it still yields sub-optimal outcomes due to the lack of a progressive harmonic branch. This demonstrates the effectiveness of harmonized attention control.
4.2.2 Ablation on Disentangled Inversion Technique
To demonstrate the soundness of Algorithm 2 and the effective disentanglement of triple branches, we can draw conclusions from Table 4: 1) Incorporating source distance into the target latent and harmonic branch, this leads to a decline in both preservation metrics and CLAP similarity score. 2) Incorporating target distance into the target branch and harmonic branch improves structural integrity and content preservation. However, this modification leads to a significant decrease in the CLAP score, indicating that while structural quality may improve, the alignment with the target prompt suffers. 3) Adding harmonic distance to the harmonic branch leads to a noticeable decline in performance. This finding showcases that introducing excessive deviations in the harmonic branch may adversely affect the overall audio quality and coherence.
@@ 4.2.2 Ablation on Disentangled Inversion Technique
$[{\bm{d}}_{src},{\bm{d}}_{src},{\bm{0}}]$ 14.28 0.17 2.64 4.67 0.57
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{src}]$ 13.17 0.17 2.55 4.59 0.58
$[{\bm{d}}_{src},{\bm{d}}_{tgt},{\bm{0}}]$ 11.24 0.15 2.44 4.32 0.56
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ 11.12 0.14 2.41 4.29 0.56
$[{\bm{d}}_{src},{\bm{d}}_{har},{\bm{0}}]$ 37.51 0.27 27.2 14.23 0.28
$[{\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 4.54 0.61
Table 4: Ablation results from the disentangled inversion technique. [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes adding the distance (line 6 in Algorithm 2). MSE means the mean square error loss between the edited audio features and source audio features.
4.3 Qualitative Results
Figure 4: Visualizations of the source audio’s mel and edited mel-spectrograms by different editing methods.
To complement our quantitative findings, we present a qualitative comparison in Figure 4. Methods such as SDEdit and inversion-based techniques often struggle to balance high editability with preserving melodic content and harmonic structure. In contrast, our Disentangled Inversion Control performs better in precise music editing while preserving structural integrity. We provide additional qualitative results in the Appendix G for further examination.
5 Conclusion
In this paper, we explored the burgeoning realm of text-guided diffusion models for audio generation, recognizing their potential to reshape source audio in alignment with specific textual prompts. We proposed the Disentangled Inversion Control to support both rigid and non-rigid editing tasks. Instead of a two-branch setting, we add an intermediate branch named the harmonic branch to progressively integrate harmonic and melodic information in music by cross-attention control and mutual self-attention control. To counteract the accumulated errors caused by DDIM inversion and CFG, we introduced a simple but effective method named disentangled inversion to separate the diffusion process into triple branches and eliminate the latent discrepancy distance in the source branch. Our comprehensive evaluations, conducted on the ZoME-Bench—a robust benchmark for music editing comprising 1,100 samples across 10 varied editing categories—attested to the superiority of our methods. And the experiment results on variable length test sets and ablation studies further validated the effectiveness and robustness of our method. We envisage that our work could serve as a basis for future zero-shot music editing studies.
Appendix A Preliminaries And Analyses
This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied to diffusion models for text-guided audio synthesis. It further explores the challenges associated with these methods.
A.1 Diffusion Models
Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition ${\bm{c}}$ to an output audio ${\bm{z}}_{0}$, corresponding to the given conditioning prompt. We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the Gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following the objective:
@@ A.1 Diffusion Models
Where noise is added to the sampled data ${\bm{z}}_{0}$ according to timestamp $t$. At inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using a pre-trained network for $T$ steps. To generate audios from given ${\bm{z}}_{T}$, we employ the deterministic DDIM sampling:
${\bm{z}}_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}{\bm{z}}_{t}+\sqrt{\alpha_{t-1}}(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ (11)
A.2 DDIM Inversion
While diffusion models have superior characteristics in the feature space that can support various down-stream tasks, it is hard to apply them to audios in the absence of natural diffusion feature space for non-generated audios. Thus, a simple inversion technique known as DDIM inversion is commonly used for unconditional diffusion models, predicated on the presumption that the ODE process can be reversed in the limit of infinitesimally small steps:
${\bm{z}}^{*}_{t}=\frac{\sqrt{\alpha_{t}}}{\sqrt{\alpha_{t-1}}}{\bm{z}}^{*}_{t-1}+\sqrt{\alpha_{t}}(\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1})\epsilon_{\theta}({\bm{z}}^{*}_{t-1},t-1)$ (12)
However, in most text-based diffusion models, this presumption cannot be guaranteed, resulting in a perturbation from ${\bm{z}}_{t}$ to ${\bm{z}}^{*}_{t}$. Consequently, an additional perturbation from ${\bm{z}}^{*}_{t}$ to ${\bm{z}}^{src}_{t}$ arises when sampling an audio from ${\bm{z}}^{*}_{T}$ where $\alpha$ is hyper-parameter:
${\bm{z}}_{t}=\sqrt{\alpha_{t}}{\bm{z}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$ (13)
A.3 Classifier-free Guidance
Classifier-free Guidance (CFG) (Ho & Salimans, 2021) is proposed to overcome the limitation of text-conditioned models, where text adherence could be weak. However, a higher guidance scale, $\omega$, which is intended to strengthen the model’s fidelity to the text prompt, inadvertently magnifies the accumulated inversion error. This becomes problematic in editing scenarios where precise control over the audio synthesis is desired. The modified noise estimation in CFG can be expressed as:
$\hat{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}},\varnothing)=\omega\cdot\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})+(1-\omega)\cdot\epsilon_{\theta}({\bm{z}}_{t},t,\varnothing)$ (14)
where $\varnothing$ = (“”) is the embedding of a null text. This further leads to another perturbation from ${\bm{z}}^{{}^{\prime}}_{t}$ to ${\bm{z}}^{{}^{\prime\prime}}$ due to the destruction of the DDIM process and causes error augmentation as demonstrated in Figure 1.
Appendix B Benchmark construction
B.1 General information
@@
Editing type id Editing type size origin prompt editing prompt editing instruction
@@
Table 5: Information of ZoME-Bench dataset
B.2 Annotation Process
We rebuild our caption from captions for Musiccaps offered by  (Agostinelli et al., 2023). With the help of ChatGPT-4 (OpenAI, 2023), we rebuild the caption with prompt as follows(take type “change melody" as examples):
Description: “There is a description of a Piece of music, Please judge whether the description has information of melody. If not, just answer “Flase", else change its melody properly into the opposite one, just change the adjective and don’t replace any instrument! ",“blended_word" is [origin melody, changed melody], “emphasize" is [changed melody], “blended_word" and “emphasize" are tuples.
@@ B.2 Annotation Process
Answer:(“source_prompt": “A mellow, [passionate] melody from a noisy electric guitar", “editing_prompt": “A mellow, [soft] melody from a noisy electric guitar", “blended_word": [“passionate melody", “soft melody"], “emphasize": [“soft melody"])
@@
Answer: (“source_prompt": “A recording of solo harp music with a dreamy, [relaxing] melody.", “editing_prompt": “A recording of solo harp music with a dreamy, [nervous] melody.", “blended_word": [“relaxing melody",“nervous melody"], “emphasize" :[“nervous melody"])
Question: (“A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions")
Answer: (“source_prompt": “A vintage, emotional song with [passionate] flute melody and soft wooden percussions.", “editing_prompt": “A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.", “blended_word": [“harmonized flute melody",“passionate flute melody"]), “emphasize" :[“passionate flute melody"])
Now we have Question:({origin caption}), Answer(?)"
In the same way, instructions are appended by prompt as follows (take type “change melody" as examples):
@@
Now we have Question: [‘source prompt’] & [‘editing prompt’], Answer(?)"

### v4

@@ 5. Experiments
Fixed Length Comparisons on ZoME-Bench
@@
[…]
Variable Length Comparisons On MusicDelta
[…]
5.1. Experimental Setup
Datasets. To address the absence of a standardized benchmark for inversion and editing techniques and to systematically validate our proposed method as a plug-and-play strategy for music editing and compare with existing zero-shot music editing methods, we construct a new benchmark ZoME-Bench (Zero-shot Music Editing Benchmark). ZoME-Bench comprises 1,100 audio samples, which are selected from MusicCaps (agostinelli2023musiclm) ZoME-Bench covers 10 different editing types that include both rigid and non-rigid modifications, detailed in supplementary material. Each sample is accompanied by its corresponding source prompt, target prompt, human instruction, and source audio.
We evaluate different editing methods using the pre-trained AudioLDM 2 model (liu2023audioldm2) with 200 inference steps, on NVIDIA A800 GPU. The setting of all baselines follows that of  manor2024zero. All baseline editing methods are first evaluated on ZoME-Bench across all editing types of 10 seconds duration. We further use the commonly used MusicDelta subset of the MedleyDB (bittner2014medleydb) dataset for variable length performance comparisons. MusicDelta comprises 34 musical excerpts in varying styles and in lengths ranging from 20 seconds to 5 minutes.
Evaluation Metrics. Our comprehensive evaluation involves both objective and subjective metrics to assess essential content preservation, text-audio alignment fidelity, and audio quality. (1) Objective Metrics: We use a variety of objective metrics to measure different aspects of audio editing effectiveness. Structure Distance (SD) (ju2023direct), CLAP Score (Contrastive Language-Audio Pretraining) (elizalde2023clap), LPAPS (Learned Perceptual Audio Patch Similarity) (iashin2021taming; paissan2023audio), Chroma (Chromagram Similarity), FAD (Fréchet Audio Distance) (kilgour2018fad), and Accuracy are included in our evaluation framework. (2) Subjective Metrics: For subjective evaluation, we conduct crowd-sourced human assessments using the Mean Opinion Score (MOS). This metric is used to evaluate both edit fidelity (MOS-Q) and content preservation (MOS-P).
5.2. Zero-shot Music Editing Results
We present a comparative study of our MEDIC with DIC against several established music generation and editing baselines. We group these baselines into generation-based methods, including AudioLDM 2 (liu2023audioldm2) and MusicGen (copet2023simple), and inversion-based methods, including SDEdit (liu2023audioldm), DDIM Inversion (ho2020denoising), MusicMagus (zhang2024musicmagus), and DDPM-Friendly (manor2024zero).
Fixed Length Comparisons. We evaluate the generated audio samples on the ZoME-Bench test set with a fixed length, focusing on the key aspects of content preservation and edit fidelity. As exhibited in Table 1, the results yield the following insights: (1) Our MEDIC substantially outperforms both generation inversion-based models in terms of edit fidelity and content preservation in both objective and subjective metrics, demonstrating its effectiveness in addressing complex editing tasks. (2) While DDPM-Friendly and MusicMagus improve content preservation (higher MOS-P), they lag in text-audio alignment and editing precision, indicated by their lower CLAP Scores and Accuracy. In contrast, MEDIC achieves consistently better CLAP, Accuracy, LPAPS, FAD, and Chroma, demonstrating superior alignment with target prompts as well as overall musical similarity.
Variable Length Comparisons We further evaluate MEDIC against the baseline methods in a variable length setting on the MusicDelta dataset. Table 1 highlights the following insights: (1) MEDIC consistently outperforms all baselines across all objective and subjective metrics, demonstrating strong robustness and adaptability for editing longer audio segments. (2) Inversion-based baselines suffer a notable drop in edit fidelity, as seen in their lower Accuracy and CLAP scores, likely due to error accumulation and insufficient attention control. In contrast, MEDIC achieves the highest CLAP and Accuracy, confirming its ability to deliver precise and well-aligned edits even on audio longer than 20 seconds.
Fine-grained Comparisons on ZoME-Bench. We further compare performance across different editing types using FAD and CLAP scores (Figure 3): (1) MEDIC consistently outperforms all baselines for both rigid and non-rigid editing tasks. (2) Baselines can handle some rigid edits, but perform poorly on non-rigid manipulations such as “Change Genre” and “Change Melody”.
Figure 6. Visualizations of the mel-spectrograms of the source audio and the edited audio by different editing methods.
5.3. Ablation Study and Visualization
Effect of Attention Control Methods. To validate our attention control designs, we conduct ablation studies on the following configurations: Remove Mutual Self-Attention Control (w/o MSA Control), Remove Cross-Attention Control (w/o CA Control), and Remove Harmonic Branch (w/o Harmonic Branch). Table 2 reveals that: (1) Both cross-attention and mutual self-attention controls significantly enhance editing performance. Their presence leads to improvements across all metrics, demonstrating their crucial role in achieving higher content preservation and edit fidelity. (2) While the naive combination of mutual self-attention and cross-attention control improves metrics such as LPAPS, FAD, and Chroma, it results in sub-optimal outcomes without the inclusion of the Harmonic Branch. This missing component reduces overall coherence and refinement, emphasizing the critical role of the Harmonic Branch in augmenting attention control mechanisms to optimize both content integrity and edit fidelity.
Table 2. Effect of Mutual Self-Attention control (MSA), Cross-Attention (CA) control, and Harmonic Branch (HB).
[…]
HAC 11.97 0.15 2.49 0.73 0.61 0.59
w/o MSA Ctrl 14.13 0.19 2.75 0.57 0.56 0.45
w/o CA Ctrl 13.78 0.18 2.50 0.63 0.58 0.48
w/o HB 12.75 0.16 2.59 0.66 0.59 0.51
Effectiveness of Disentangled Inversion Technique. We assess the soundness of Algorithm 2 and the effectiveness of disentanglement of triple branches. As shown in Table 3, 1) When employing ${\bm{d}}_{src},{\bm{0}},{\bm{0}}]$—the configuration used in MEDIC—results are notably strong across several metrics. MEDIC achieves the highest Accuracy (0.59), CLAP score (0.61), and Chroma (0.73), demonstrating its balanced approach in effectively preserving content fidelity while maintaining strong alignment with the target prompt. 2) Configurations that incorporate additional distances, such as ${\bm{d}}_{src},{\bm{d}}_{src},{\bm{0}}]$ and ${\bm{d}}_{src},{\bm{d}}_{har},{\bm{0}}]$, show a notable decline in performance across most metrics. 3)Interestingly, the configuration ${\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ achieves the lowest values for LPAPS and FAD, suggesting potential improvements in audio similarity. However, this setup seems to markedly reduce overall text-audio alignment and accuracy, indicating a trade-off where structural precision comes at the expense of editorial coherence.
Table 3. Ablation study of Disentangled Inversion Technique. $\cdot$$\cdot,\cdot$] denotes adding the distance (line 6 in Algorithm 2). MSE is the mean square error loss between the edited audio features and source audio features. ${\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ is used in MEDIC.
[…]
${\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 0.73 0.61 0.59
${\bm{d}}_{src},{\bm{d}}_{src},{\bm{0}}]$ 14.28 0.17 2.64 0.53 0.57 0.47
${\bm{d}}_{src},{\bm{0}},{\bm{d}}_{src}]$ 13.17 0.17 2.55 0.47 0.58 0.52
${\bm{d}}_{src},{\bm{d}}_{tgt},{\bm{0}}]$ 11.24 0.15 2.44 0.52 0.56 0.52
${\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ 11.12 0.14 2.41 0.55 0.56 0.32
${\bm{d}}_{src},{\bm{d}}_{har},{\bm{0}}]$ 37.51 0.27 2.72 0.32 0.28 0.31
${\bm{0}},{\bm{0}},{\bm{0}}]$ 12.65 0.16 2.54 0.47 0.57 0.47
Visualization. To complement our quantitative findings, we present a qualitative comparison in Figure 6. Methods such as SDEdit and inversion-based techniques often struggle to balance high editability and preserve melodic content and harmonic structure. In contrast, MEDIC performs better in precise music editing while preserving structural integrity. We provide additional qualitative results for all editing categories in supplementary material, demonstrating the superiority of our approach.
6. Conclusion
In this paper, we propose the Disentangled Inversion Control to support both rigid and non-rigid editing tasks and develop a zero-shot music editing framework MelodyEdit based on DIC. We add an intermediate harmonic branch to progressively integrate harmonic and melodic information in music by cross-attention control and mutual self-attention control. To counteract the accumulated errors caused by DDIM inversion and CFG, we introduce Disentangled Inversion to separate the diffusion process into triple branches and eliminate the latent discrepancy distance in the source branch. Extensive experiments demonstrate superiority of MelodyEdit on both fixed and variable length settings. We envisage that our work could serve as a basis for future zero-shot music editing studies.
Acknowledgements.
The research was supported by the Early Career Scheme (ECS-Hong Kong University of Science and Technology 22201322) and NSFC (No. 62206234) of Mainland China.
  1. (1)
    Benchmark construction
    1. (a)
      General information
    2. (b)
      Annotation Process
    3. (c)
      Data Format
  2. (2)
    Implementation Details
    1. (a)
      Metrics
  3. (3)
    Quantitative Results
  4. (4)
    Potential Negative Societal Impacts
  5. (5)
    Limitations
  6. (6)
    Qualitative Results
    1. (a)
      Change Instrument
    2. (b)
      Add Instrument
    3. (c)
      Delete Instrument
    4. (d)
      Change Genre
    5. (e)
      Change Mood
    6. (f)
      Change Rhythm
    7. (g)
      Change Background
    8. (h)
      Change Melody
    9. (i)
      Extract Instrument
  7. (7)
    safeguards
7. Benchmark construction
7.1. General information
@@
Editing
type id
Editing type size origin prompt editing prompt editing instruction
@@
Table 4. Information of ZoME-Bench dataset
7.2. Annotation Process
We rebuild our caption from captions for Musiccaps offered by  (agostinelli2023musiclm). With the help of ChatGPT-4 (openai2023gpt4), we rebuild the caption with prompt as follows(take type “change melody” as examples):
Description: “There is a description of a Piece of music, Please judge whether the description has information of melody. If not, just answer “Flase”, else change its melody properly into the opposite one, just change the adjective and don’t replace any instrument! ”,“blended_word” is [origin melody, changed melody], “emphasize” is [changed melody], “blended_word” and “emphasize” are tuples.
@@ 7.2. Annotation Process
Answer:(“source_prompt”: “A mellow, [passionate] melody from a noisy electric guitar”, “editing_prompt”: “A mellow, [soft] melody from a noisy electric guitar”, “blended_word”: [“passionate melody”, “soft melody”], “emphasize”: [“soft melody”])
@@
Answer: (“source_prompt”: “A recording of solo harp music with a dreamy, [relaxing] melody.”, “editing_prompt”: “A recording of solo harp music with a dreamy, [nervous] melody.”, “blended_word”: [“relaxing melody”,“nervous melody”], “emphasize” :[“nervous melody”])
Question: (“A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions”)
Answer: (“source_prompt”: “A vintage, emotional song with [passionate] flute melody and soft wooden percussions.”, “editing_prompt”: “A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.”, “blended_word”: [“harmonized flute melody”,“passionate flute melody”]), “emphasize” :[“passionate flute melody”])
Now we have Question:({origin caption}), Answer(?)”
In the same way, instructions are appended by prompt as follows (take type “change melody” as examples):
@@
Now we have Question: [‘source prompt’] & [‘editing prompt’], Answer(?)”